OpenStudy (chrisneedshelp):

Given the system of equations, what is the solution? 2x + y = -1 x - y = -5

OpenStudy (chrisneedshelp):

@d3Xter

OpenStudy (anonymous):

Given: \(\begin{align}2x + y = -1 \\x - y = -5 \end{align}\) Add both eqns : \(3x = -6 \longrightarrow x = \frac{-6}{3} = -2\)

OpenStudy (anonymous):

Subst \(x = -2 \) in 2nd eqn to get y \(\Longrightarrow \begin{align}-2 -y = -5 \\ -y = 2-5 \\y=3\end{align}\)

OpenStudy (chrisneedshelp):

{(-8, -3)} {(2, 7)} {(-2, 3)}